Solve the linear equation 2x+3y=1, 5x+7y=3
Answer by waynest(281)

2x + 3y = 1
5x + 7y = 3
change the first equation by multiplying it by 5; and change the second by multiplying it by 2
10x + 15y = 5
10x + 14y = 6
subtract the second equation from the first giving:
y = -1
replace y in the first equation with -1
2x + 3(-1) = 1
2x - 3 = 1
2x -3 + 3 = 1 + 3
2x = 4
2x/2 = 4/2
x = 2
check:
2(2) + 3(-1) = 1
4 - 3 = 1
1 = 1
5(2) + 7(-1) = 3
10 - 7 = 3
3 = 3